THE MEASUREMENT OF AMBIENT NO_{2} BY CAVITY RING-DOWN SPECTROSCOPY

\underline{JAMES HARGROVE}, LIMING WANG, AIKO MATSANUGA, SUSAN ZAIDE, and JINGSONG ZHANG, Department of Chemistry and Air Pollution Research Center, University of California, Riverside, CA 92521; KAREN MUYSKENS and MARK MUYSKENS, Department of Chemistry, Calvin College, Grand Rapids, MI 49546.

Nitrogen dioxide (NO_{2}) is a principle component of photochemical smog that is present in polluted atmospheres at the parts per billion (ppb) level. In this work, cavity ring-down spectroscopy (CRDS) is utilized for the detection of ambient NO_{2}. Sub ppb limit of detection is routinely achieved. Experimental approaches to remove the particulates and to improve the detection limit and accuracy are investigated. Ambient measurements of NO_{2} by CRDS are evaluated, and are compared side by side with the NO_{2} measurements by chemiluminescence analysis.